# Circuit-tuning: A Mechanistic Approach for Identifying Parameter Redundancy and Fine-tuning Neural Networks

### v1

@@ C.3 Analysis of the Quality of the Discovered Circuit
Figure 4: The influence from top n edges.
C.3.1 Faithfulness and Completeness
Faithfulness and completeness examine a circuit from two different views. Faithfulness tells how much performance a circuit gets, while completeness tells how much performance a circuit fails to capture. Consider a model $M$ with its computational graph $\mathcal{G}$, a circuit $\mathcal{C}$ for a specific task $T$ and a metric $\mathcal{L}_{m}$ for measuring the output of the model, following the definition in (Marks et al., 2024), the faithfulness of the circuit $\mathcal{C}$ is
$\frac{\mathcal{L}_{m}[M(\mathcal{C})]-\mathcal{L}_{m}[M(\varnothing)]}{\mathcal{L}_{m}[M(\mathcal{G})]-\mathcal{L}_{m}[M(\mathcal{\varnothing})]}$
@@ C.3.1 Faithfulness and Completeness
$\frac{\mathcal{L}_{m}[M(\mathcal{G}\backslash\mathcal{C})]-\mathcal{L}_{m}[M(\varnothing)]}{\mathcal{L}_{m}[M(\mathcal{G})]-\mathcal{L}_{m}[M(\mathcal{\varnothing})]}$
@@
C.3.2 Ablation Study of Random activation
@@
C.4 Analyses of the Training Dynamics
C.4.1 Interpret the Circuit for the Subject-verb Disagreement Task
@@
Subject Identification Heads To find out the Subject Identification Heads, we check the attention pattern of each attention head at the END token since the END token is directly used for predicting the verb token. We sort the heads in descending order according to their attention weights from the END token (query) to the subject tokens (key). Then we keep the heads in which the attention is mainly paid to the subject in the context.
We find that there exist two types of Subject Identification Heads. The type I heads mainly attend to the last token itself at the END token, so the attention pattern is a diagonal line. This type of head is helpful when the subject is exactly the END token, e.g. “He is …”, “The girls are …”, etc. The type II heads attend to the subject which is several tokens before the END token, e.g. “The kid who is holding an ice cream in hand is …”, “The famous scientist, who is also an artist, has …”, etc. The type II heads obviously have the ability of syntactic analysis and subject identification, while the type I heads may just happen to attend to the subject that overlaps with the last token.
@@
Figure 7: The examples of the attention patterns in the Subject Attribute Heads. In Figure 7(a), the END token “video” attends to the subject “video” which is also the END token itself. In Figure 7(b), the END token “opponents” mainly attends to the subject “interaction”.
@@
Subject Attribute Heads To find out the Subject Attribute Heads, we need to find out which heads directly affect the match between the verb and the subject, which is measured by the logit difference between the flipped verb and the original verb at the END token. Suppose the output of the final layer at the END token is $x_{END}\in\mathbb{R}^{D}$, then the logit difference is $[W_{U}(v_{flip})-W_{U}(v))]\cdot x_{END}$, in which $W_{U}(v_{flip})-W_{U}(v)$ is called the logit lens (nostalgebraist, 2020). Since the logit lens is fixed, we expect the projection of $x_{END}$ on the direction of the logit lens to be large, thus encouraging the probability difference between the two opposite verb forms (love v.s. loves, etc.). As discussed in Appendix B, the output $x_{END}$ which is in the residual stream can be decomposed into the linear addition of the outputs from the previous layers. Therefore, the output of a Subject Attribute Head is a part of $x_{END}$ and would encourage the value of the logit difference. Thus, a Subject Attribute Head is an attention head that has a large dot product value with the logit lens.
@@
In practice, we calculate the dot product between the logit lens $W_{U}(v_{flip})-W_{U}(v)$ and the output $Attn_{i}^{j}(x)$ from each attention head in each layer over a batch of samples. The result is shown in the main text in Figure 2. The darker the color, the larger the absolute value of the dot product is, which implies that the head is more likely to be a Subject Attribute Head. We observe that head.6.0, head.6.5, head.8.5, head.10.9, and so on see obvious flips (Figure 8) from positive to negative or the opposite direction, which implies that they are directly responsible for the match between the subject and the verb. During training, the parameters inside these heads adjust themselves to the new data distribution, while their function type remains unchanged, which is an interesting finding of the self-regulation ability inside the model.
Figure 8: A sketch of the flip from subject-verb agreement to disagreement.
@@
Collaborative Heads Finally, we notice that some of the Subject Attribute Heads (head.8.5, head.10.9, etc.) are also Subject Identification Heads, which means they also attend to subject tokens. We wonder if there exist some heads in previous layers that could influence the behavior of the Subject Attribute Heads. That is to say, the Subject Attribute Heads do not act alone but collaborate with other heads.
@@
To find out these heads, we knock out the upstream heads one at a time at the END token using mean ablation. We observe the change in the attention pattern of each Subject Attribute Head and then keep the heads that bring obvious changes in the attention patterns. In practice, we focus on subject-verb agreement and only check the influence on head.8.5 and head.10.9. We also provide two types of data, corresponding to the two cases discussed in Section C.4.1, in order to provide a more detailed analysis. Results show that
@@
C.4.2 The Interaction and Collaboration Inside the Model
@@
    Each node is responsible for a single or multiple functions. As discussed in Section C.4.1, we have found attention heads that are responsible for identifying the subject in a sentence or adjust the form of a verb according to the attributes of the subject, or both. Each head has its division of labor when completing a task.
@@
C.4.3 The Evidence of Hebbian Learning
As discussed in Section 5.1.5, we observe that some edges in the circuit are strengthened or weakened during training, just like the Hebbian learning proposed in (Do, 1949) in neuroscience. As stated by Hebb, a synapse between two neurons is strengthened when the neurons on either side of the synapse have highly correlated outputs, which means they are often activated synchronously. The theory is often concluded as “Cells that fire together, wire together” (Shatz, 1992). For two neurons $i$ and $j$, a common description of hebbian learning is as follows:
$w_{ij}=\frac{1}{p}\sum_{k=1}^{p}x_{i}^{k}x_{j}^{k}$
@@ C.4.3 The Evidence of Hebbian Learning
We also find that the edge contribution may keep decreasing during training. This is quite similar to the self-organization of cells inside human brains, a well-known phenomenon of which is the lateral inhibition among neurons (Jacobson, 1993), which means an activated neuron can reduce the activity of its neighbors. Inspired by this, (Kohonen, 2012) developed the self-organizing maps (SOM), an unsupervised algorithm that leverages the Winner-Take-All strategy to perform competitive learning. The core ideas behind SOM are:
@@
In our study, we find that the dynamic change of edges echoes the above discussions on competition and self-organization. During training, some connections between nodes are strengthened, which may reduce the intensity of other connections. After training, the components inside a model have reorganized themselves to adapt to the new data distribution. When faced with an input, different regions inside the computational graph are responsible for different subtasks and collaborate to complete a goal, as discussed in Section C.4.1 and Section C.4.2.
Table 3: Some of the strengthened and weakened edges during training. The dynamic change shows the change in edge contribution $log[1+c(e)]$ (the start of training $\rightarrow$ 2000 steps $\rightarrow$ 3000 steps $\rightarrow$ 4000 steps). The dynamic process is visualized in Figure 9.
Strengthened Edges Weakened Edges
edge dynamic change edge dynamic change
mlp.2 $\rightarrow$ mlp.3 0 $\rightarrow$ 0.0293 $\rightarrow$ 0.0993 $\rightarrow$ 0.2282 mlp.0 $\rightarrow$ mlp.10 0.2712 $\rightarrow$ 0.058 $\rightarrow$ 0 $\rightarrow$ 0
(a) circuit at 2000 steps
(b) circuit at 3000 steps
(c) circuit at 4000 steps
Figure 9: Evidence of Hebbian learning in the subject-verb disagreement task
C.4.4 The Circuits Before and After Fine-tuning
We present the circuits of the subject-verb disagreement task before and after fine-tuning in Figure 10(a) and Figure 10(b) respectively. The circuit of subject-verb disagreement is trained under the setting of $N=1000$ edges, and for both of the circuits we only present the top 100 edges. It can be seen that most of the heads we discussed in Section C.4.1 are included in the circuit. Besides, it is obvious that the circuits before and after fine-tuning share a lot of nodes, which implies that the function shift from subject-verb agreement to disagreement happens mostly in the polarity change of nodes, instead of randomly assigning the ability to other nodes. This is an intuitive finding, which not only demonstrates the rationality of circuit-tuning as well as our analyses but also provides new insights for our understanding of the mechanism inside language models.
@@ C.4.4 The Circuits Before and After Fine-tuning
C.4.5 The Experiment Results after the Revision of Attribution Score
During the analyses in Section C.4.1, we find that the original definition of the attribution score in EAP (Syed et al., 2023) fails to capture all the relevant edges in a task. For example, head.6.0 which is a Subject Attribute Head fails to appear in the circuit. As discussed in Section 5.1.6, we assume that there exists a situation where an important node is connected with many other nodes, but each edge is not that strong. For example, as illustrated in Figure 11, an upstream node $n_{a}^{u}$ is connected with four downstream nodes, while another upstream node $n_{b}^{u}$ is connected with only one downstream node. Since the edge between $n_{b}^{u}$ and $n_{5}^{d}$ is stronger than any edge between $n_{a}^{u}$ and the nodes connected with it, the edge $n_{b}^{u}\rightarrow n_{5}^{d}$ may be kept in the circuit, while the edges in $\mathcal{E}_{a}=\{n_{a}^{u}\rightarrow n_{i}^{d}|i=1,2,3,4\}$ may be left out. As a result, $n_{a}^{u}$ is not involved in optimization, though the sum of the edge contributions of all edges in $\mathcal{E}_{a}$ may be almost the same or even larger than that of $n_{b}^{u}\rightarrow n_{5}^{d}$.
@@ C.4.5 The Experiment Results after the Revision of Attribution Score
$c(e)^{\prime}=c(e)\cdot\sum_{k=1}^{N_{down}^{i}}c(n_{i}^{u}\rightarrow n_{k}^{d})\cdot\sum_{k=1}^{N_{up}^{j}}c(n_{k}^{u}\rightarrow n_{j}^{d})$
where $c(e)$ is the original attribution score in EAP, $N_{down}^{i}$ is the number of the downstream nodes of $n_{i}^{u}$, $N_{up}^{j}$ is the number of the upstream nodes of $n_{i}^{u}$. The revision considers the contributions from all the edges connected to the upstream node and the downstream node. To verify it, we conduct a new experiment on the subject-verb disagreement task and compare it with the result before. Results are shown in Figure 1(c). Details can be found in Table 4. Compared with the original attribution score, our method improves the logit difference steadily, while even bringing down the computation to some extent.
@@
Appendix D Details for the Complex Tasks
D.1 Details for Task Settings
D.1.1 Reasoning-based Tasks
Mathematics We use GSM8K (Cobbe et al., 2021) as the dataset, which contains about 8.5k grade school math problems with natural language solutions. The answer to a problem not only contains the final answer but also provides the process for solving the problem. An example of this task is shown in Table 5.
@@ D.1.1 Reasoning-based Tasks
Logical Reasoning We use Contexthub (Hua et al., 2024) as the dataset, which consists of problems of 4 difficulty levels, including deductive and abductive reasoning in 12 distinct categories or domains from Wikipedia. The problems together with their reasoning processes are instantiated automatically by LLMs following the fixed formal logic templates. The whole dataset contains 18,240 samples. We only use level 1 and level 2 in our experiment for convenience, which include 6720 samples in total. An example of this task is shown in Table 6.
@@
Reasoning (aaa OR aab) $\rightarrow$ aac = False. Given aac is False, the value of premise (aaa OR aab) is False, thus, the value of aab is abduced as False. Thus, the answer is False “The Sahara desert receives heavy rainfall this year” or “The Amazon rainforest experiences severe drought conditions”) logically implies “Some of Earth’s major ecosystems are undergoing significant changes in weather patterns” whose corresponding truth value is False. Given “Some of Earth’s major ecosystems are undergoing significant changes in weather patterns” is False, the value of premise (”The Sahara desert receives heavy rainfall this year” or “The Amazon rainforest experiences severe drought conditions”) is False, thus, the value of “The Amazon rainforest experiences severe drought conditionsïs abduced as False. Thus, the answer is <answer>False</answer>
@@
For all reasoning-based tasks, we add an instruction “Please answer step by step.” at the end of the question in order to guide the model to answer the question step by step.
D.1.2 Reasoning-free Tasks
Gender De-biasing According to (Gallegos et al., 2024), there are various kinds of expressions in social bias. In this study, we focus on the gender bias in occupations. We aims to break down the binary gender stereotype of a model. For example, given a sentence “the doctor put on [PRP] coat” where [PRP] is a possessive pronoun, we expect the model to choose his or her with equal probabilities.
During fine-tuning, the model learns to predict the next word in an auto-regressive way, thus we expect the model to balance the probabilities between male attribute words (he/his/him/himself) and female attribute words (she/her/herself) at the END token when generating the next token. Therefore, we use the logit difference between the male attribute words and female attribute words at the END token as the metric $\mathcal{L}_{m}$ for measuring the output of the model. Specifically, for each sample, we calculate the logit difference between the pronoun and the anti-pronoun, which is the pronoun in the opposite gender. For example, in the case “the doctor put on [PRP] coat”, the logit difference is $logit(W_{her}|W_{END})-logit(W_{his}|W_{END})$, where $W_{END}$ is the END token on.
We use BUG (Levy et al., 2021) for training, which is a large-scale dataset of sentences sampled from real-world corpora. Each sentence is marked with an occupation and the pronouns referring to it. In practice, we use the “balanced BUG” provided in the dataset which includes 2.5w sentences randomly sampled from BUG to ensure balance between male and female entities and between stereotypical and non-stereotypical gender role assignments. We perform coreference resolution ourselves to filter out the samples in which the coreference is not right, and leave 1.5w samples for training, which contains 151 types of occupations and each sentence only contains one (occupation, pronoun) pair.

### v2

@@ F.3 Analysis of the Quality of the Discovered Circuit
Figure 7: The influence from the setting of top $N$ edges.
F.3.1 Faithfulness and Completeness
Faithfulness and completeness examine a circuit from two different views. Faithfulness tells how much performance a circuit gets, while completeness tells how much performance a circuit fails to capture. Consider a model $M$ with its computational graph $\mathcal{G}$, a circuit $\mathcal{C}$ for a specific task $T$ and a metric $\mathcal{L}_{m}$ for measuring the output of the model, following the definition in [34], the faithfulness of the circuit $\mathcal{C}$ is
$\frac{\mathcal{L}_{m}[M(\mathcal{C})]-\mathcal{L}_{m}[M(\varnothing)]}{\mathcal{L}_{m}[M(\mathcal{G})]-\mathcal{L}_{m}[M(\mathcal{\varnothing})]}$ (28)
@@ F.3.1 Faithfulness and Completeness
$\frac{\mathcal{L}_{m}[M(\mathcal{G}\backslash\mathcal{C})]-\mathcal{L}_{m}[M(\varnothing)]}{\mathcal{L}_{m}[M(\mathcal{G})]-\mathcal{L}_{m}[M(\mathcal{\varnothing})]}$ (29)
@@
F.3.2 Ablation Study of Random activation
@@
F.4 Analyses of the Training Dynamics
F.4.1 Interpret the Circuit for the Subject-verb Disagreement Task
@@
Subject Identification Heads
To find out the Subject Identification Heads, we check the attention pattern of each attention head at the END token since the END token is directly used for predicting the verb token. We sort the heads in descending order according to their attention weights from the END token (query) to the subject tokens (key). Then we keep the heads in which the attention is mainly paid to the subject in the context.
We find that there exist two types of Subject Identification Heads. The type I heads mainly attend to the last token itself at the END token, so the attention pattern is a diagonal line. This type of head is helpful when the subject is exactly the END token, e.g. “He is …", “The girls are …", etc. The type II heads attend to the subject which is several tokens before the END token, e.g. “The kid who is holding an ice cream in hand is …", “The famous scientist, who is also an artist, has …", etc. The type II heads obviously have the ability of syntactic analysis and subject identification, while the type I heads may just happen to attend to the subject that overlaps with the last token.
@@
Figure 10: The examples of the attention patterns in the Subject Attribute Heads. In Figure 10(a), the END token “video" attends to the subject “video" which is also the END token itself. In Figure 10(b), the END token “opponents" mainly attends to the subject “interaction".
Subject Attribute Heads
@@
To find out the Subject Attribute Heads, we need to find out which heads directly affect the match between the verb and the subject, which is measured by the logit difference between the flipped verb and the original verb at the END token. Suppose the output of the final layer at the END token is $x_{END}\in\mathbb{R}^{D}$, then the logit difference is $[W_{U}(v_{flip})-W_{U}(v))]\cdot x_{END}$, in which $W_{U}(v_{flip})-W_{U}(v)$ is called the logit lens [40]. Since the logit lens is fixed, we expect the projection of $x_{END}$ on the direction of the logit lens to be large, thus encouraging the probability difference between the two opposite verb forms (love v.s. loves, etc.). As discussed in Appendix C, the output $x_{END}$ which is in the residual stream can be decomposed into the linear addition of the outputs from the previous layers. Therefore, the output of a Subject Attribute Head is a part of $x_{END}$ and would encourage the value of the logit difference. Thus, a Subject Attribute Head is an attention head that has a large dot product value with the logit lens.
@@
In practice, we calculate the dot product between the logit lens $W_{U}(v_{flip})-W_{U}(v)$ and the output $Attn_{i}^{j}(x)$ from each attention head in each layer over a batch of samples. The result is shown in the main text in Figure 3. The darker the color, the larger the absolute value of the dot product is, which implies that the head is more likely to be a Subject Attribute Head. We observe that head.6.0, head.6.5, head.8.5, head.10.9, and so on see obvious flips (Figure 3(a)) from positive to negative or the opposite direction, which implies that they are directly responsible for the match between the subject and the verb. During training, the parameters inside these heads adjust themselves to the new data distribution, while their function type remains unchanged, which is an interesting finding of the self-regulation ability inside the model.
Collaborative Heads
@@
Finally, we notice that some of the Subject Attribute Heads (head.8.5, head.10.9, etc.) are also Subject Identification Heads, which means they also attend to subject tokens. We wonder if there exist some heads in previous layers that could influence the behavior of the Subject Attribute Heads. That is to say, the Subject Attribute Heads do not act alone but collaborate with other heads.
@@
To find out these heads, we knock out the upstream heads one at a time at the END token using mean ablation. We observe the change in the attention pattern of each Subject Attribute Head and then keep the heads that bring obvious changes in the attention patterns. In practice, we focus on subject-verb agreement and only check the influence on head.8.5 and head.10.9. We also provide two types of data, corresponding to the two cases discussed in Appendix F.4.1, in order to provide a more detailed analysis. Results show that
@@
F.4.2 The Interaction and Collaboration Inside the Model
@@
    Each node is responsible for a single or multiple functions. As discussed in Appendix F.4.1, we have found attention heads that are responsible for identifying the subject in a sentence or adjust the form of a verb according to the attributes of the subject, or both. Each head has its division of labor when completing a task.
@@
F.4.3 The Evidence of Hebbian Learning
As discussed in Section 5.1.5, we observe that some edges in the circuit are strengthened or weakened during training, just like the Hebbian learning proposed in [10] in neuroscience. As stated by Hebb, a synapse between two neurons is strengthened when the neurons on either side of the synapse have highly correlated outputs, which means they are often activated synchronously. The theory is often concluded as “Cells that fire together, wire together" [54]. For two neurons $i$ and $j$, a common description of hebbian learning is as follows:
$w_{ij}=\frac{1}{p}\sum_{k=1}^{p}x_{i}^{k}x_{j}^{k}$ (30)
@@ F.4.3 The Evidence of Hebbian Learning
We also find that the edge contribution may keep decreasing during training. This is quite similar to the self-organization of cells inside human brains, a well-known phenomenon of which is the lateral inhibition among neurons [22], which means an activated neuron can reduce the activity of its neighbors. Inspired by this, [25] developed the self-organizing maps (SOM), an unsupervised algorithm that leverages the Winner-Take-All strategy to perform competitive learning. The core ideas behind SOM are:
@@
In our study, we find that the dynamic change of edges echoes the above discussions on competition and self-organization. During training, some connections between nodes are strengthened, which may reduce the intensity of other connections. After training, the components inside a model have reorganized themselves to adapt to the new data distribution. When faced with an input, different regions inside the computational graph are responsible for different subtasks and collaborate to complete a goal, as discussed in Appendix F.4.1 and Appendix F.4.2.
F.4.4 The Circuits Before and After Fine-tuning
We present the circuits of the subject-verb disagreement task before and after fine-tuning in Figure 11(a) and Figure 11(b) respectively. The circuit of subject-verb disagreement is trained under the setting of $N=1000$ edges, and for both of the circuits we only present the top 100 edges. It can be seen that most of the heads we discussed in Appendix F.4.1 are included in the circuit. Besides, it is obvious that the circuits before and after fine-tuning share a lot of nodes, which implies that the function shift from subject-verb agreement to disagreement happens mostly in the polarity change of nodes, instead of randomly assigning the ability to other nodes. This is an intuitive finding, which not only demonstrates the rationality of circuit-tuning as well as our analyses but also provides new insights for our understanding of the mechanism inside language models.
@@ F.4.4 The Circuits Before and After Fine-tuning
F.4.5 The Experiment Results after the Revision of Attribution Score
During the analyses in Appendix F.4.1, we find that the original definition of the attribution score in EAP [56] fails to capture all the relevant edges in a task. For example, head.6.0 which is a Subject Attribute Head fails to appear in the circuit. We assume that there exists a situation where an important node is connected with many other nodes, but each edge is not that strong. For example, as illustrated in Figure 12, an upstream node $n_{a}^{u}$ is connected with four downstream nodes, while another upstream node $n_{b}^{u}$ is connected with only one downstream node. Since the edge between $n_{b}^{u}$ and $n_{5}^{d}$ is stronger than any edge between $n_{a}^{u}$ and the nodes connected with it, the edge $n_{b}^{u}\rightarrow n_{5}^{d}$ may be kept in the circuit, while the edges in $\mathcal{E}_{a}=\{n_{a}^{u}\rightarrow n_{i}^{d}|i=1,2,3,4\}$ may be left out. As a result, $n_{a}^{u}$ is not involved in optimization, though the sum of the edge contributions of all edges in $\mathcal{E}_{a}$ may be almost the same or even larger than that of $n_{b}^{u}\rightarrow n_{5}^{d}$.
@@ F.4.5 The Experiment Results after the Revision of Attribution Score
$c(e)^{\prime}=c(e)\cdot\sum_{k=1}^{N_{down}^{i}}c(n_{i}^{u}\rightarrow n_{k}^{d})\cdot\sum_{k=1}^{N_{up}^{j}}c(n_{k}^{u}\rightarrow n_{j}^{d})$ (31)
where $c(e)$ is the original attribution score in EAP, $N_{down}^{i}$ is the number of the downstream nodes of $n_{i}^{u}$, $N_{up}^{j}$ is the number of the upstream nodes of $n_{i}^{u}$. The revision considers the contributions from all the edges connected to the upstream node and the downstream node. To verify it, we conduct a new experiment on the subject-verb disagreement task and compare it with the result before. Results are shown in 13. Details can be found in Table 4. Compared with the original attribution score, our method improves the logit difference steadily, while even bringing down the computation to some extent.
@@
Figure 13: The experiment result after the revision on attribution score.
@@
Appendix G Details for the Complex Tasks
G.1 Details for Task Settings
G.1.1 Reasoning-based Tasks
Mathematics We use GSM8K [8] as the dataset, which contains about 8.5k grade school math problems with natural language solutions. The answer to a problem not only contains the final answer but also provides the process for solving the problem. An example of this task is shown in Table 5.
@@ G.1.1 Reasoning-based Tasks
Logical Reasoning We use Contexthub [21] as the dataset, which consists of problems of 4 difficulty levels, including deductive and abductive reasoning in 12 distinct categories or domains from Wikipedia. The problems together with their reasoning processes are instantiated automatically by LLMs following the fixed formal logic templates. The whole dataset contains 18,240 samples. We only use level 1 and level 2 in our experiment for convenience, which include 6720 samples in total. An example of this task is shown in Table 6.
@@
Reasoning (aaa OR aab) $\rightarrow$ aac = False. Given aac is False, the value of premise (aaa OR aab) is False, thus, the value of aab is abduced as False. Thus, the answer is False “The Sahara desert receives heavy rainfall this year" or “The Amazon rainforest experiences severe drought conditions") logically implies “Some of Earth’s major ecosystems are undergoing significant changes in weather patterns" whose corresponding truth value is False. Given “Some of Earth’s major ecosystems are undergoing significant changes in weather patterns" is False, the value of premise ("The Sahara desert receives heavy rainfall this year" or “The Amazon rainforest experiences severe drought conditions") is False, thus, the value of “The Amazon rainforest experiences severe drought conditionsïs abduced as False. Thus, the answer is <answer>False</answer>
@@
For all reasoning-based tasks, we add an instruction “Please answer step by step." at the end of the question in order to guide the model to answer the question step by step.
G.1.2 Reasoning-free Tasks
Gender De-biasing According to [16], there are various kinds of expressions in social bias. In this study, we focus on the gender bias in occupations. We aims to break down the binary gender stereotype of a model. For example, given a sentence “the doctor put on [PRP] coat" where [PRP] is a possessive pronoun, we expect the model to choose his or her with equal probabilities.
During fine-tuning, the model learns to predict the next word in an auto-regressive way, thus we expect the model to balance the probabilities between male attribute words (he/his/him/himself) and female attribute words (she/her/herself) at the END token when generating the next token. Therefore, we use the logit difference between the male attribute words and female attribute words at the END token as the metric $\mathcal{L}_{m}$ for measuring the output of the model. Specifically, for each sample, we calculate the logit difference between the pronoun and the anti-pronoun, which is the pronoun in the opposite gender. For example, in the case “the doctor put on [PRP] coat", the logit difference is $logit(W_{her}|W_{END})-logit(W_{his}|W_{END})$, where $W_{END}$ is the END token on.
We use BUG [30] for training, which is a large-scale dataset of sentences sampled from real-world corpora. Each sentence is marked with an occupation and the pronouns referring to it. In practice, we use the “balanced BUG" provided in the dataset which includes 2.5w sentences randomly sampled from BUG to ensure balance between male and female entities and between stereotypical and non-stereotypical gender role assignments. We perform coreference resolution ourselves to filter out the samples in which the coreference is not right, and leave 1.5w samples for training, which contains 151 types of occupations and each sentence only contains one (occupation, pronoun) pair.